OpenStudy (anonymous):

Rae earns $60 for playing music on Saturday nights at the local coffee house. She makes an average of $5.25 in tips per hour. Write the function of Rae's earnings, and solve for how much she would make after 4 hours.

OpenStudy (anonymous):

Okay so it looks to me like we can plug this into y=mx+b and solve.

OpenStudy (anonymous):

So 60 can be the y-intercept because we already know she is for sure getting that much. And 5.25 can be m. The hours will act as x.

OpenStudy (anonymous):

\[y=5.25(4)+60\]

OpenStudy (ccswims):

function is f(x) = mx + b not y = mx + b It doesn't look like a big difference but it is

OpenStudy (ccswims):

its the difference between a function and a linear equation

OpenStudy (anonymous):

@ccswims lol no it really does not matter. But yes when writing out the function put f(x)=5.25(4)+60. I do not think points would have been taken off though

OpenStudy (anonymous):

Anyways, 5.25 multiplied by 4 is 21. Just add 21 and 60 and that is how much money she earned(:

OpenStudy (ccswims):

idk, my teacher used to take off points in tests if we did y = instead of f(x) = when we were still in the Functions Chapter